OpenStudy (soahelp):

Very confused please help. A line passes through (3, 7) and (6, 9). Which equation best represents the line? y = 3 over 2x + 5 y = 2 over 3x + 5 y = 3x + 2 y = 2 over 3x + 2

OpenStudy (anonymous):

Can you find slope for the points given to you?

OpenStudy (anonymous):

Slope is given by: \[Slope = \frac{y_2 - y_1}{x_2 -x_1}\]

OpenStudy (soahelp):

4/3

OpenStudy (soahelp):

3/4

OpenStudy (tkhunny):

9-7 = ?? 6-3 = ??

OpenStudy (soahelp):

9-7=3 6-3= 3

OpenStudy (anonymous):

oh, wow..!! \(9-7 = 3\) ?? Are you sure?

OpenStudy (soahelp):

no its two

OpenStudy (anonymous):

Good..

OpenStudy (soahelp):

and i found the answer to my problem

OpenStudy (anonymous):

That is nice.. :)

OpenStudy (soahelp):

thxs for your help anyway

OpenStudy (soahelp):

its b

OpenStudy (anonymous):

So, what is the answer you got?

OpenStudy (anonymous):

How can you say it is not D) ?

OpenStudy (soahelp):

because when i added the two i did not get my answer

OpenStudy (anonymous):

Why you added two?